OpenStudy (anonymous):

Rewrite 22 plus 33 using GCF and the distributive property

OpenStudy (judygreeneyes):

22 = 2 * 11 33 = 3 * 11 22 + 33 = 11(2 + 3) (I think this is what you want...)